A drawer contains 4 red socks, 6 white socks, and 10 blue socks. Without looking, you draw out a sock and then draw out a second sock without returning the first sock. What is the probability that the first sock and the second sock are both white.
Answer by stanbon(48569)

A drawer contains 4 red socks, 6 white socks, and 10 blue socks. Without looking, you draw out a sock and then draw out a second sock without returning the first sock. What is the probability that the first sock and the second sock are both white.
--------
# of ways to draw two white socks: 6C2=15
# of ways to draw two socks of any color: 20C2=190
Prob(draw two white socks) = 15/190 = 3/38
------------
Another way to do it:
Prob draw 1st white sock: 6/20
Prob draw 2nd white sock when 1st is not replaced = 5/19
Prob(two white) = 6/20 *5/19 = 3/10 * 5/19 = 3/38
==========
Cheers,
stan H.